Use the following information for questions a) – f). Bank Assets Risk Weight Bank Liabilities Cash...

  1. Use the following information for questions a) – f).

Bank Assets

Risk Weight

Bank Liabilities

Cash and Treasury Securities

$8,000

0%

Deposits

$32,000

Municipal Bonds

$6,000

20%

Hot Money

$25,000

Repurchase Agreements

$4,000

20%

Subordinated Debt

$1,000

CMOs

$10,000

50%

Common Stock

$1,000

Single Family Home Mortgage

$15,000

50%

Surplus

$1,200

Commercial Loans

$10,000

100%

Retained Earnings

$1,800

Agriculture Loans

$5,000

100%

Allowance for Loan Loss

($3,000)

0%

Bank Buildings

7,000

100%

Total

$62,000

$62,000

  1. How much Tier 1 capital does the bank have?
  2. What is the minimum leverage capital for this bank?
  3. What is the amount of risk-adjusted assets for the bank?
  4. What is the minimum Tier 1 capital for this bank is?
  5. What is the minimum total capital for this bank?
  6. What is the total amount of the bank's regulatory capital?
Answer #1

Soultion:

1. How much Tier 1 capital does the bank have?

Tier 1 capital includes common shares, stock surpluses, common shares issued by subsidiaries, retained earnings and accumulated OCI.

Tier 1 capital = Common stock + Surplus + Retained earnings = 1000 + 1200 + 1800 = $4,000.

2. What is the minimum leverage capital for this bank?

Minimum leverage capital = 3% of bank total assets = 0.03 * 62000 = $1,860

3. What is the amount of risk-adjusted assets for the bank?

Risk adjusted assets = 0*8000 + 0.2*6000 + 0.2*4000 + 0.5*10000 + 0.5*5000 + 1*10000 + 1*5000 + 0*-3000 + 1*7000 = $36,500

4. What is the minimum Tier 1 capital for this bank is?

Minimum Tier 1 capital = 6% of bank risk-adjusted assets = 0.06 * 36500 = $2,190.

What is the minimum total capital for this bank?

Minimum Total capital = 8% of bank risk-adjusted assets = 0.08 * 36500 = $2,920.

5. What is the total amount of the bank's regulatory capital?

Total amount of the bank's regulatory capital = Tier 1 capital + Tier 2 capital

Tier 1 capital = Common stock + Surplus + Retained earnings = 1000 + 1200 + 1800 = $4,000.

Tier 2 capital = Subordinated debt = $1,000

Total amount of the bank's regulatory capital = 4000 + 1000 = $5,000